# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v2

@@ 3.1 Relaxed control formulation
  • The space $D^{n}[0,T]$ is endowed with the Skorokhod metric $d_{D}$ and the corresponding Borel $\sigma$-algebra denoted by $\mathcal{F}^{D}$. Let $\mathcal{F}_{t}^{D}$ be the Borel $\sigma$-algebra up to time $t$ (c.f. [20]) for $t\in[0,T]$.
  • The space $\mathcal{Q}$ of relaxed controls is defined as the set of measures $q$ in $0,T]\times U$ with the first marginal equal to the Lebesgue measure and $\int_{[0,T]\times U}|u|^{p}q(\mathrm{d}t,\mathrm{d}u)<\infty$. We endow the space $\mathcal{Q}$ with the $2$-Wasserstein metric on $\mathcal{P}_{2}([0,T]\times U)$ given by $d_{\mathcal{Q}}(q^{1},q^{2})=\mathcal{W}_{2,[0,T]\times U}\left(\frac{q^{1}}{T},\frac{q^{2}}{T}\right)$, where the metric on $0,T]\times U$ is given by $((t_{1},u_{1}),(t_{2},u_{2}))\mapsto|t_{2}-t_{1}|+|u_{2}-u_{1}|$. Note that each $q\in\mathcal{Q}$ can be identified with a measurable function $0,T]\in t\mapsto q_{t}\in\mathcal{P}_{2}(U)$, defined uniquely up to $\mathrm{a.s.}$ by $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$. In the sequel, we will always refer to the measurable mapping $q=(q_{t})_{t\in[0,T]}$ to a relaxed control in $\mathcal{Q}$. Let $\mathcal{F}^{\mathcal{Q}}$ be the Borel $\sigma$-algebra of $\mathcal{Q}$ and $\mathcal{F}_{t}^{\mathcal{Q}}$ be the $\sigma$-algebra generated by the maps $q\mapsto q([0,s]\times V)$ with $s\in[0,t]$ and Borel measurable $V\subset U$.
  • The space $A^{k}[0,T]$ is a subset of $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$, where the set $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$ is the collection of mappings $a:[0,T]\to\mathbb{R}^{k}$ such that $a$ is càdlàg and $a(0)=0$. As in [20], there is a natural one-to-one correspondence between $\mathcal{V}$ and $(\mathcal{M}[0,T])^{k}$, that is, for $a=(a^{1},\ldots,a^{k})\in\mathcal{V}$ and $i=1,\ldots,k$, define $\nu_{i}((s,t])=a^{i}(t)-a^{i}(s)$ for $0\leq s\leq t\leq T$. Then, $\mathcal{V}$ is the dual space of $(C[0,T])^{k}$, and hence we endow $\mathcal{V}$ with the corresponding weak topology (which is also equivalent to the Fortet-Mourier norm on ($\mathcal{M}[0,T])^{k}$, c.f. [5]). Thus, $A^{k}[0,T]$ is a closed subset of $(\mathcal{M}[0,T])^{k}$ and can also be viewed as $(\mathcal{M}^{+}[0,T])^{k}$. Let $\mathcal{F}^{A}$ be the Borel $\sigma$-algebra and $\mathcal{F}^{A}_{t}$ be the Borel $\sigma$-algebra defined up to time $t\in[0,T]$ (cf. [20]).
@@
$\Omega:=\widehat{\Omega}\times A^{k}[0,T]=D^{n}[0,T]\times\mathcal{Q}\times A^{k}[0,T].$ (4)
[…]
@@
    for all $g\in C_{b}^{2}(\mathbb{R}^{n})$, $M^{R}g=(M_{t}^{R}g)_{t\in[0,T]}$ is an $(R,\mathbb{F})$-martingale, where
    $M_{t}^{R}g$ $:=g(X_{t})-\int_{0}^{t}\int_{U}\mathbb{G}g(s,X_{s},R^{X}(s),u)q_{s}(\mathrm{d}u)\mathrm{d}s-\int_{0}^{t}\nabla_{x}g(X_{s-})G(s)\mathrm{d}\zeta_{s}$
    $\quad-\sum_{0<s\leq t}\left(g(X_{s})-g(X_{s-})-\nabla_{x}g(X_{s-})\Delta X_{s}\right)$
@@
    $\mathbb{G}g(t,x,\mu,u):=b(t,x,\mu,u)\nabla_{x}g(x)+\frac{1}{2}\mathrm{tr}\left(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(x)\right).$
@@
    for $i\in I$, $m\otimes R(A_{i}^{R})=T$ with the admissible set $A_{i}^{R}$ defined by
@@
    where $m\otimes R$ denotes the product measure of Lebesugue measure $m$ and $R$.
@@ 3.2 Existence of optimal relaxed controls
For $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$, let $\mathcal{R}_{\rm ad}(\nu)$ be the set of admissible control rules $R$ in Definition 3.1. For $R\in\mathcal{R}_{\rm ad}(\nu)$, we have the following martingagle measure characterization:
@@
$R\in\mathcal{R}_{\rm ad}(\nu)$ iff there is a filtered probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime}=(\mathcal{F}^{\prime}_{t})_{t\in[0,T]},P^{\prime})$ supporting an $n$-dimensional $\mathbb{F}^{\prime}$-adapted process $X=(X_{t})_{t\in[0,T]}$, $r$-dimensional orthogonal $\mathbb{F}^{\prime}$-martingale measures $N=(N^{1},\ldots,N^{r})$ on $U\times[0,T]$, each with $\mathcal{F}^{\prime}_{t}$-progressively measurable intensity $q_{t}(\mathrm{d}u)\mathrm{d}t$ and a $k$-dimensional $\mathbb{F}^{\prime}$-progressively measurable nondecreasing process $\zeta=(\zeta_{t})_{t\in[0,T]}$ such that $R=P^{\prime}\circ(X,q,\zeta)^{-1}$, and the following conditions hold:
@@
$\Gamma(R):=\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right].$ (6)
@@
$\mathcal{R}_{\rm opt}(\nu):=\inf\{R\in\mathcal{R}_{\rm ad}(\nu);~\Gamma(R)=v^{*}\}.$ (7)
@@
Motivated by this expression, we have the next result whose proof is omitted because the necessary part is obvious and the sufficiency can be shown by the classical martingale argument.
@@
    $\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$; (iv) for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, $\tilde{M}_{t}^{R}g$ is a $(R,\mathbb{F})$-martingale, where $\tilde{M}_{t}^{R}g:=g(Y_{t})-\int_{0}^{t}\int_{U}\tilde{\mathbb{G}}g(s,X_{s},R^{X}(s),u)(Y_{s})q_{s}(\mathrm{d}u)\mathrm{d}s$ with
    $\tilde{\mathbb{G}}g(t,x,\mu,u)(y):=b(t,x,\mu,u)\nabla_{x}g(y)+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}g(y)).$
@@
$\Lambda(R,\omega)=\int_{0}^{T}\int_{U}f(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}(\omega).$
@@
$v^{*}+\epsilon$ $\geq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right]$
$\geq\mathbb{E}^{R}\left[-\int_{0}^{T}M(1+|X_{t}|^{p}+M_{2}(\mu))\mathrm{d}t+M_{1}\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+c_{0}|\zeta_{T}|\right].$
@@
Then, it follows from Corollary 2.9 in [20] that $(\tilde{R}_{\ell}|_{\mathcal{A}^{k}[0,T]})_{\ell\geq 1}$ is tight. Moreover, by Proposition B.3 in Lacker [24] and Kolmogorov’s tightness criterion, we can deduce that $(\tilde{R}_{\ell}|_{\mathcal{Q}\times C([0,T];\mathbb{R}^{n})})_{\ell\geq 1}$ is relatively compact in $\mathcal{P}_{2}(\mathcal{Q}\times C([0,T];\mathbb{R}^{n}))$.
@@
which proves claim (i) of Lemma 3.5. For the martingale argument, because $(b,\sigma)$ are uniformly continuous in $\mu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ (actually there exists a uniform Lipschitz constant) w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, for any $g\in C_{b}^{2}(\mathbb{R}^{n})$, it holds that
$\tilde{M}_{t}^{R_{\ell}}g(\omega)\to\tilde{M}_{t}^{R}g(\omega),\quad\ell\to\infty$ (11)
@@
$\left|\mathbb{E}^{R}\left[(\tilde{M}_{t}^{R}g(\omega)-\tilde{M}_{s}^{R}g(\omega))\eta(\omega)\right]\right|\leq\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{t}^{R}g(\omega)-\tilde{M}_{t}^{R_{n}}g(\omega)\right)\eta(\omega)\right|\right]$
$\quad+\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{s}^{R}g(\omega)-\tilde{M}_{s}^{R_{\ell}}g(\omega)\right)\eta(\omega)\right|\right]+\left|\mathbb{E}^{R}\left[\tilde{M}_{t}^{R}g(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{t}^{R}g(\omega)\eta(\omega)\right]\right|$
$\quad+\left|\mathbb{E}^{R}\left[\tilde{M}_{s}^{R}g(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{s}^{R}g(\omega)\eta(\omega)\right]\right|+\left|\mathbb{E}^{R_{\ell}}\left[\left(\tilde{M}_{t}^{R_{\ell}}g(\omega)-\tilde{M}_{s}^{R_{\ell}}g(\omega)\right)\eta(\omega)\right]\right|$

### v3

@@ 3.1 Relaxed control formulation
  • The space $D^{n}[0,T]$ is endowed with the Skorokhod metric $d_{D}$ and the corresponding Borel $\sigma$-algebra denoted by $\mathcal{F}^{D}$. Let $\mathcal{F}_{t}^{D}$ be the naturally generated Borel $\sigma$-algebra up to time $t\in[0,T]$ (c.f. [25]).
  • The space $\mathcal{Q}$ of relaxed controls is defined as the set of measures $q$ in $0,T]\times U$ with the first marginal equal to the Lebesgue measure and $\int_{[0,T]\times U}|u|^{p}q(\mathrm{d}t,\mathrm{d}u)<\infty$. We endow the space $\mathcal{Q}$ with the $2$-Wasserstein metric on $\mathcal{P}_{2}([0,T]\times U)$ given by $d_{\mathcal{Q}}(q^{1},q^{2})=\mathcal{W}_{2,[0,T]\times U}\left(\frac{q^{1}}{T},\frac{q^{2}}{T}\right)$, where the metric on $0,T]\times U$ is given by $((t_{1},u_{1}),(t_{2},u_{2}))\mapsto|t_{2}-t_{1}|+|u_{2}-u_{1}|$. Note that each $q\in\mathcal{Q}$ can be identified with a measurable function $0,T]\in t\mapsto q_{t}\in\mathcal{P}_{2}(U)$, defined uniquely up to $\mathrm{a.s.}$ by $q(\mathrm{d}t,\mathrm{d}u)=q_{t}(\mathrm{d}u)\mathrm{d}t$. In the sequel, we will always refer to the measurable mapping $q=(q_{t})_{t\in[0,T]}$ to a relaxed control in $\mathcal{Q}$. Let $\mathcal{F}^{\mathcal{Q}}$ be the Borel $\sigma$-algebra of $\mathcal{Q}$ and $\mathcal{F}_{t}^{\mathcal{Q}}$ be the $\sigma$-algebra generated by the maps $q\mapsto q([0,s]\times V)$ with $s\in[0,t]$ and Borel measurable $V\subset U$. By the same argument of Lemma 3.2 in [29], one can verify that $q=(q_{t})_{t\in[0,T]}$ is $\mathbb{F}^{\mathcal{Q}}=(\mathcal{F}_{t}^{\mathcal{Q}})_{t\in[0,T]}$-progressively measurable.
  • The space $A^{k}[0,T]$ is a subset of $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$, where the set $\mathcal{V}\subset\mathrm{BV}^{k}[0,T]$ is the collection of mappings $a:[0,T]\to\mathbb{R}^{k}$ such that $a$ is càdlàg and $a(0)\geq 0$. As in [25] and Section 3.5 of [21], there is a natural one-to-one correspondence between $\mathcal{V}$ and $(\mathcal{M}[0,T])^{k}$, that is, for $a=(a^{1},\ldots,a^{k})\in\mathcal{V}$ and $i=1,\ldots,k$, define $\nu_{i}(\{0\})=a_{0}^{i}$ and $\nu_{i}((s,t])=a^{i}(t)-a^{i}(s)$ for $0\leq s\leq t\leq T$. Then, $\mathcal{V}$ is the dual space of $(C[0,T])^{k}$, and hence we endow $\mathcal{V}$ with the corresponding weak topology (which is also equivalent to the Fortet-Mourier norm on ($\mathcal{M}[0,T])^{k}$ introduced in (5) below, or c.f. [6]). Thus, $A^{k}[0,T]$ is a closed subset of $(\mathcal{M}[0,T])^{k}$ by Portmaneau theorem ($\nu^{i}(\{0\})\geq\limsup_{n\to\infty}\nu_{i}^{n}(\{0\})\geq 0$ for any $i=1,\dots,k$) and can also be viewed as $(\mathcal{M}^{+}[0,T])^{k}$. Let $\mathcal{F}^{A}$ be the Borel $\sigma$-algebra and $\mathcal{F}^{A}_{t}$ be the naturally generated Borel $\sigma$-algebra defined up to time $t\in[0,T]$ (cf. [25]).
@@
$\Omega:=\widehat{\Omega}\times A^{k}[0,T],\quad{\widehat{\Omega}=D^{n}[0,T]\times\mathcal{Q}.}$ (4)
Equip $\widehat{\Omega}$ and $\Omega$ with the respective product metrics $d_{\widehat{\Omega}}(\widehat{\omega}^{1},\widehat{\omega}^{2})=d_{D}(\psi^{1},\psi^{2})+d_{\mathcal{Q}}(q^{1},q^{2})$ and $d_{\Omega}(\omega^{1},\omega^{2})=d_{\widehat{\Omega}}(\widehat{\omega}^{1},\widehat{\omega}^{2})+\sum_{i=1}^{k}d_{\rm FM}(\nu_{i}^{1},\nu_{i}^{2})$ for $\omega^{i}=(\widehat{\omega}^{i},\zeta^{i})=(\psi^{i},q^{i},\zeta^{i})\in\Omega$ with $i=1,2$. Here, $\nu^{i}$ is the correspondence of $\zeta^{i}$ in $\mathcal{V}$ and $d_{\rm FM}$ is the Fortet-Mourier metric on $\mathcal{M}[0,T]$ that, for any $\mu,\nu\in{\cal M}[0,T]$,
$d_{\rm FM}(\mu,\nu):=\sup_{\begin{subarray}{c}\|f\|_{\infty}+{\rm Lip}(f)\leq 1\end{subarray}}\left|\int_{[0,T]}f\,d\mu-\int_{[0,T]}f\,d\nu\right|,$ (5)
which induces the weak convergence of finite measures. We adopt this metric as it is compatible with weak convergence while preserving completeness and separability, ensuring that the resulting product spaces remain Polish.
@@
[…]
$d_{\rm FM,\Omega}(\mu,\nu):=\sup_{\begin{subarray}{c}\|f\|_{\infty}\leq 1+{\rm Lip}(f)\leq 1\end{subarray}}\left|\int_{\Omega}f\,d\mu-\int_{\Omega}f\,d\nu\right|.$ (6)
@@
    for all $\psi\in C_{b}^{2}(\mathbb{R}^{n})$, $M^{R}\psi=(M_{t}^{R}\psi)_{t\in[0,T]}$ is an $(R,\mathbb{F})$-martingale, where
    $M_{t}^{R}\psi$ $:=\psi(X_{t})-\int_{0}^{t}\int_{U}\mathbb{G}\psi(s,X_{s},R^{X}(s),u)q_{s}(\mathrm{d}u)\mathrm{d}s-\int_{0}^{t}\nabla_{x}\psi(X_{s-})G(s)\mathrm{d}\zeta_{s}$
    $\quad-\sum_{0<s\leq t}\left(\psi(X_{s})-\psi(X_{s-})-\nabla_{x}\psi(X_{s-})\Delta X_{s}\right)$
@@
    $\mathbb{G}\psi(t,x,\mu,u):=b(t,x,\mu,u)\nabla_{x}\psi(x)+\frac{1}{2}\mathrm{tr}\left(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}\psi(x)\right).$
@@
    for $i\in I$, $(m\times R)(A_{i}^{R})=T$ with the admissible set $A_{i}^{R}$ defined by
@@
    where $m\otimes R$ denotes the product measure of Lebesgue measure $m$ and $R$.
Remark 3.2.
To the best of our knowledge, the formulation of the relaxed constraint condition in Definition 3.1-(iv) is novel within the compactification approach. Unlike the existing work, we incorporate the original state-control-law constraint directly into the definition of admissible probability measures on the canonical space. Such definition, to the best of our knowledge, has not been considered before in the compactification literature.
@@ 3.2 Existence of optimal relaxed controls
For $\nu\in\mathcal{P}_{p}(\mathbb{R}^{n})$, let $\mathcal{R}_{\rm ad}(\nu)$ be the set of admissible control rules $R$ in Definition 3.1. For $R\in\mathcal{R}_{\rm ad}(\nu)$, we have the following martingale measure characterization:
@@
$R\in\mathcal{R}_{\rm ad}(\nu)$ iff there is a filtered probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{F}^{\prime}=(\mathcal{F}^{\prime}_{t})_{t\in[0,T]},P^{\prime})$ supporting an $n$-dimensional $\mathbb{F}^{\prime}$-adapted process $X=(X_{t})_{t\in[0,T]}$, $r$-dimensional orthogonal $\mathbb{F}^{\prime}$-martingale measures $N=(N^{1},\ldots,N^{r})$ on $U\times[0,T]$, each with $\mathcal{F}^{\prime}_{t}$-progressively measurable intensity $q_{t}(\mathrm{d}u)\mathrm{d}t$ and a $k$-dimensional $\mathbb{F}^{\prime}$-progressively measurable non-decreasing process $\zeta=(\zeta_{t})_{t\in[0,T]}$ such that $R=P^{\prime}\circ(X,q,\zeta)^{-1}$, and the following conditions hold:
@@
$\Gamma(R):=\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+{\int_{[0,T]}c(t)d\zeta_{t}+g(X_{T},R^{X}(T))}\right].$ (8)
@@
$\mathcal{R}_{\rm opt}(\nu):={\{R\in\mathcal{R}_{\rm ad}(\nu):~\Gamma(R)=v^{*}\}}.$ (9)
@@
Motivated by this expression, we have the next result whose proof is omitted because the necessary part is obvious and the sufficiency can be shown by the classical martingale argument (c.f. Proposition 3.1 in [25]).
@@
    $\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+|\zeta_{T}|^{p}\right]<\infty$; (iv) for any $\psi\in C_{b}^{2}(\mathbb{R}^{n})$, $\tilde{M}_{t}^{R}\psi$ is a $(R,\mathbb{F})$-martingale, where $\tilde{M}_{t}^{R}\psi:=\psi(Y_{t})-\int_{0}^{t}\int_{U}\tilde{\mathbb{G}}\psi(s,X_{s},R^{X}(s),u)(Y_{s})q_{s}(\mathrm{d}u)\mathrm{d}s$ with
    $\tilde{\mathbb{G}}\psi(t,x,\mu,u)(y):=b(t,x,\mu,u)\nabla_{x}\psi(y)+\frac{1}{2}\mathrm{tr}(\sigma\sigma^{\top}(t,x,\mu,u)\nabla_{xx}^{2}\psi(y)).$
@@
$\Lambda(R,\omega)=\int_{0}^{T}\int_{U}f(t,X_{t}(\omega),R^{X}(t),u)q_{t}(\omega,\mathrm{d}u)\mathrm{d}t+{\int_{[0,T]}c(t)d\zeta_{t}(\omega)+g(X_{T}(\omega),R^{X}(T))}.$
@@
$v^{*}+\epsilon$ $\geq\mathbb{E}^{R}\left[\int_{0}^{T}\int_{U}f(t,X_{t},R^{X}(t),u)q_{t}(\mathrm{d}u)\mathrm{d}t+{\int_{[0,T]}c(t)d\zeta_{t}+g(X_{T},R^{X}(T))}\right]$
$\geq\mathbb{E}^{R}\Big[-\int_{0}^{T}M(1+|X_{t}|^{p}+M_{2}(R^{X}(t)))\mathrm{d}t+M_{1}\int_{0}^{T}\int_{U}|u|^{p}q_{t}(\mathrm{d}u)\mathrm{d}t+c_{0}|\zeta_{T}|$
$\qquad\quad-M(1+|X_{T}|^{p}+M_{2}(R^{X}(T)))\Big].$
@@
Then, it follows from Corollary 2.9 in [25] that $(\tilde{R}_{\ell}|_{\mathcal{A}^{k}[0,T]})_{\ell\geq 1}$ is tight. Moreover, by Proposition B.3 in [29] and Kolmogorov’s tightness criterion, we can deduce that $(\tilde{R}_{\ell}|_{\mathcal{Q}\times C([0,T];\mathbb{R}^{n})})_{\ell\geq 1}$ is relatively compact in $\mathcal{P}_{2}(\mathcal{Q}\times C([0,T];\mathbb{R}^{n}))$.
@@
which proves claim (i) of Lemma 3.6. For the martingale argument, because $(b,\sigma)$ are uniformly continuous in $\mu\in\mathcal{P}_{p}(\mathbb{R}^{n})$ (actually there exists a uniform Lipschitz constant) w.r.t. $(t,x,u)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}$, for any $\psi\in C_{b}^{2}(\mathbb{R}^{n})$, it holds that
$\tilde{M}_{t}^{R_{\ell}}\psi(\omega)\to\tilde{M}_{t}^{R}\psi(\omega),\quad\ell\to\infty$ (13)
@@
$\left|\mathbb{E}^{R}\left[(\tilde{M}_{t}^{R}\psi(\omega)-\tilde{M}_{s}^{R}\psi(\omega))\eta(\omega)\right]\right|\leq\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{t}^{R}\psi(\omega)-\tilde{M}_{t}^{R_{n}}\psi(\omega)\right)\eta(\omega)\right|\right]$
$\quad+\mathbb{E}^{R_{\ell}}\left[\left|\left(\tilde{M}_{s}^{R}\psi(\omega)-\tilde{M}_{s}^{R_{\ell}}\psi(\omega)\right)\eta(\omega)\right|\right]+\left|\mathbb{E}^{R}\left[\tilde{M}_{t}^{R}\psi(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{t}^{R}\psi(\omega)\eta(\omega)\right]\right|$
$\quad+\left|\mathbb{E}^{R}\left[\tilde{M}_{s}^{R}\psi(\omega)\eta(\omega)\right]-\mathbb{E}^{R_{\ell}}\left[\tilde{M}_{s}^{R}\psi(\omega)\eta(\omega)\right]\right|+\left|\mathbb{E}^{R_{\ell}}\left[\left(\tilde{M}_{t}^{R_{\ell}}\psi(\omega)-\tilde{M}_{s}^{R_{\ell}}\psi(\omega)\right)\eta(\omega)\right]\right|$